OpenStudy (anonymous):

1 quick math help question

OpenStudy (anonymous):

a man throws a ball off the top of a building and records the height of the ball at different times as shown in the table: Height of the ball time (s) Height (feet) 0 46 1 63 2 48 3 1 a. Find a quadratic model for the data I know the answer for a is y=-16x^2+33x+46

OpenStudy (mathmath333):

let the time be \(\large t\) hence \(\Large f(t)=ax^2+bx+c\) \(\large\tt \begin{align} \color{black}{ f(t)=ax^2+bx+c\\~\\ f(0)=46\\~\\ f(t)=ax^2+bx+c\\~\\ f(0)=a(0)^2+b(0)+c\\~\\ 46=c\\~\\ c=46------\color{red}{ 1}\\~\\ f(1)=63\\~\\ f(1)=a(1)^2+b(1)+c\\~\\ 6=a+b+c\\~\\ a+b+c=6------\color{red}{ 2}\\~\\ f(2)=48\\~\\ f(2)=a(2)^2+b(2)+c\\~\\ 48=4a+2b+c------\color{red}{ 3}\\~\\\\~\\ \\~\\}\end{align}\)

OpenStudy (mathmath333):

from eqn 1 ,2 and 3 u have to find \(\Large a\) and \(\Large b\) , we have found \(\Large c\) then put the value of \(\Large a\) , \(\Large b\) and \(\Large c\) in \(\Large f(t)=ax^2+bx+c \\ \implies \Large y=ax^2+bx+c \\ \)

OpenStudy (mathmath333):

* correction a+b+c=63

OpenStudy (anonymous):

Thank you so much! That helped a bunch!

OpenStudy (mathmath333):

oh welcome